OpenStudy (anonymous):

WILL FAN AND GIVE MEDAL In the formula below, represents the electric power of a light bulb measured in watts, is the voltage, and is the current in amperes. A 90-watt light bulb has a current of 6 amperes and a voltage of 15 volts. If the voltage of the 90-watt light bulb doubles, what will happen to the current? A. The current will decrease by half. B. The current will double. C. The current will decrease by one-fourth. D. The current will stay the same.

OpenStudy (johnweldon1993):

Okay since we know that power is \[\large P = VI\] we had \[\large P = 15*6 = 90W\] Now, if we double the voltage (15V) of the light bulb, we will have \[\large P = VI\] \[\large 90 = 30*A\] what does A equal?

OpenStudy (anonymous):

3?

OpenStudy (johnweldon1993):

Right So if the Current went from 6 amps to 3 amps....

OpenStudy (anonymous):

The current will decrease by half so the answer will be A. Thanks

OpenStudy (johnweldon1993):

No problem!

OpenStudy (anonymous):

how come i can't fan you?